### v1

@@ 4.1 Technology-based Drivers
Utilizing hosting services offers an efficient way to share code, data, and ML model parameter settings, thus supporting the reproducibility of ML-driven research [71]. Examples of hosting services include the runtime environments of ML platforms. If the original author runs the ML experiment in such a runtime environment, e.g., Kaggle Kernels [^3] https://www.kaggle.com/, Google Colab [^4] https://colab.google/, or CodaLab [^5] https://codalab.org/, researchers attempting to reproduce the results should be able to execute the experiment within the same environment.
@@
Reproducing the environment and setup of any ML experiment requires the consideration of existing dependencies and software versions, and is usually a complex task itself. Virtualization can simplify this process by bundling the essential components of ML models and experiments, such as the environment and code, into a single package for being shared with other researchers. Thus, if the authors of a paper build the experiment in a virtual environment, issues associated with setup reproduction can be greatly reduced. However, the adoption of virtualization by researchers depends on its user-friendliness and the effort of integration into their current workflows [11]. Concerns about virtualization include its limitations in allowing researchers to build upon them in a scalable manner. Traditional virtual machines (VMs) emulate an entire operating system for setting up and running experiments. The use of containerization software like Docker [^6] https://www.docker.com/ has just recently become popular. Containers are more lightweight and flexible than VMs, making it easier to adapt environments for follow-up studies [11]. There are also designated platforms for computational research, such as Code Ocean [^7] https://codeocean.com/, that offer virtualization via so-called reproducible capsules. Their focus in particular is to simplify the virtualization process and allow researchers to focus on the research questions instead of the standardization of environments [15]. Additionally, there are many other tools, such as e.g., ReproZip [14], which is one of the recommended tools by the SIGMOD Reproducibility Committee to streamline reproducibility, and DetTrace [51], which aims to ensure completely deterministic computations.
@@
Many different sources of randomness during ML training lead to irreproducibility of ML models. Managing these sources of randomness, e.g., via random number seeds, deterministic algorithms, or other methods, could therefore greatly increase reproducibility. Random number seeds should be used and published to make ML experiments more reproducible and control a number of sources of the inherent nondeterminism. A seed is a first value used to initialize the pseudo-random number generator. When the same seed is used, the sequence of pseudo-random numbers generated is deterministic, meaning it will be the same every time the code is run. Additionally, ML models should be benchmarked and evaluated with multiple random number seeds, such that the variance can be reported and inform about the true performance of an ML model [63]. However, fixing the random number seeds can only eliminate algorithmic sources of randomness. Any differences due to implementation-level variance between low-level libraries could be overcome by implementing deterministic code, which has been demonstrated before [43]. Furthermore, hardware nondeterminism is also avoidable by using patching, i.e., replacing non-deterministic operations from hardware with deterministic ones, however, this will lead to a significant computing overhead [13]. To overcome the obstacle of inherent nondeterminism in Reinforcement Learning (RL), there are frameworks, which provide reproducible environments and are commonly used in RL research [12, 21, 78]. These frameworks act as benchmarking environments where different RL algorithms can be evaluated in a common environment. Additionally, different papers have outlined the implementation of deterministic RL algorithms, by controlling all the sources of non-determinism [50].
@@
Privacy-preserving technologies support reproducibility, as they enable the collaborative training of ML models without sharing private or sensitive data. The main benefit of this is, that ML models can make use of larger, and more diverse data, thus helping to decrease bias and leading to more reproducible ML models. The main aim of Privacy-Preserving Machine Learning (PPML) is to facilitate the use of privacy-sensitive data to create better ML models, and, to allow data owners to collaboratively train ML models on private data. In that regard, PPML has several requirements. First, protecting the confidentiality of the training data. Second, preventing the leakage of sensitive information from ML model parameters and outputs, i.e., to hinder the re-identification of individuals. Third, achieving the listed security aspects while still preserving the utility of the ML model [76]. To achieve this, a number of different techniques are being used and developed, such as differential privacy, fully homomorphic encryption, secure multi-party computation, and federated learning [19]. These techniques are implemented in software libraries such as TensorFlow Privacy, PySyft, ML Privacy Meter, CryptFlow, or Crypten [7]. Furthermore, data can be made anonymous by removing identifiable personal information. However, if too much data is removed, the ML models may perform poorly. If not enough data is removed, it may still be possible to re-identify individuals by combining many different non-unique features [76]. An alternative approach to PPML is to generate synthetic data that captures the same information as the original data. A robust technique for creating such datasets can produce readily available datasets of nearly any size, as demonstrated in Biomedical fields. This approach has led to the development of Synthea, a software package designed to generate synthetic patients and electronic health care records [73].
@@
There are many tools and platforms that assist in the implementation and management of ML model and ML-based applications ranging from . Some of them provide functionalities to increase ML reproducibility. A recent study has evaluated 19 ML tools to gain insight into their concepts constituting reproducibility support [61]. As a result, five main pillars of ML reproducibility in tools and platforms were identified: (i) code versioning, (ii) data access, (iii) data versioning, (iv) experiment logging, and (v) pipeline creation. Most of these pillars are associated with managing and keeping track of different artifacts created during phases of the ML lifecycle (i.e., design, development, and deployment) as for instance, datasets, labels, code, logs, environment dependencies, random number seeds or hyperparameters [67]. Each of these artifacts influences the final results of the ML model. Consequently, most tools aim to collect, store, and manage these artifacts, ensuring researchers can access and use them during reproduction attempts. Notable are also various tools and platforms for experiment tracking [67]:
  • DVC [^8] https://dvc.org/ is a version control system for ML projects with a command-line interface similar to Git [^9] https://git-scm.com/. It integrates with Git, supports cloud storage, and handles large dataset versioning. DVC ensures full code and data provenance by enabling experiment tracking.
  • MLflow [^10] https://mlflow.org/ is an open-source tool for supporting ML experiment tracking, ML model deployment, and centralized model storage. Additionally, it provides an easy-to-use user Web dashboard.
@@
  • dToolAI [31] collects and packages ML models together with supplemental information, such as hyperparameters, appropriate metadata, and persistent URIs for model training data. In contrast to the other tools, dToolAI is specifically tailored for deep learning models.
AutoML platforms, such as H2O Driverless AI [^11] https://h2o.ai, Google Cloud AutoML [^12] https://cloud.google.com/automl, DataRobot [^13] https://www.datarobot.com/platform) are a novel subcategory of ML tools that aim to aid with every aspect of the ML lifecycle, from data aggregation to model deployment. AutoML tools could thus facilitate more standardized ML models and also take care of tasks like hyperparameter optimization. Furthermore, many of them also provide a lot of the features identified as the main pillars of reproducibility in ML tools. It is however questionable how practical these tools are for research, since they often act as black-boxes without the need to apply programming expertise. Sometimes, e.g., in the case of H2O Driverless AI, the features go a step further and promise to be able to prevent problems such as overfitting or target leakage. In the case of leakage, this is done by checking for a strong correlation between a feature and the target and then taking action, e.g., warning the user or automatically handling it. This is, however, a very simple solution to the problem and does not address the more complex cases of data leakage that are often present in research, e.g., temporal leakage. [^14] https://h2o.ai/wiki/target-leakage//
@@ 4.2 Procedural Drivers
Due to a lack of shared datasets, currently, many researchers in ML-driven research - most notably in Biomedical fields - use individually acquired data [45]. The collection of such data is a time-consuming task and bears a significant risk of causing reproducibility issues, e.g., bias or data leakage. Often, the data population of these datasets is not very large and, thus, might not translate well to new contexts. Creating shared and standardized datasets can, therefore, (i) save researchers time in acquiring new data, (ii) facilitate the collaborative and independent maintenance and verification of data to minimize methodological errors, and (iii) support transferability and generalizability, through the use of multi-institutional data [45]. In addition to the standardization of datasets, data cards [60] provide a consistent and comparable framework for reporting essential aspects of ML datasets. This includes information e.g., about access restrictions, risks and limitations associated with the usage of the dataset, or any preprocessing steps, amongst many other contents, which are needed for responsible ML development.
Another issue is the lack of standardized evaluation, which leads to reported performances of ML models often being overly optimistic [18]. To ensure the statistical significance of ML model evaluations, it is crucial to report performance as an aggregate of results obtained from multiple random runs with different random number seeds [16]. Furthermore, ML models should, if possible, be tested and evaluated on multiple different datasets [20]. This underscores the need for standardized evaluation methods, which can be supported by checklists or tools to prevent errors in this critical aspect of ML research. For this, similarly to data cards, model cards [46] are aimed to standardize the evaluation and reporting of the performance of ML models for a variety of use cases. Model cards should inform users about the possible applications of the ML model and its limitations. For the creation of the model cards, a toolkit is also being developed by Google [^15] https://research.google/blog/introducing-the-model-card-toolkit-for-easier-model-transparency-reporting/. In RL, the creation of standardized evaluation pipelines is continually being researched to enable reproducible benchmarking of different RL algorithms [40].
@@
There are many guidelines and checklists that outline best practices for increasing the reproducibility of ML. The guidelines are often aimed at specific parts of the ML workflow. The FAIR guiding principles [^16] https://www.go-fair.org/fair-principles/, e.g., aim to improve the management and stewardship of scientific data by making scientific data (i) findable, (ii) accessible, (iii) interoperable, and (iv) reusable. Other guidelines promote the transparency and openness of scientific reporting in general, such as the TOP guidelines [^17] https://www.cos.io/initiatives/top-guidelines, which target journals. Similarly, checklists provide a simple framework for ensuring certain criteria are met. Checklists have been applied effectively in the past, e.g., in safety-critical systems, where they were used as early as in 1935 to complete pre-flight checks in Boeing airplanes. A promising example is the ML checklist proposed in [57]. It has been mentioned during the NeurIPS 2019 reproducibility challenge, and has been suggested as best practice by researchers of different fields, e.g., in chemistry [6]. The checklist requests information about (1) the models and algorithms being used, (2) theoretical claims in the research article, (3) data, (4) code, and (5) the ML experiment(s). However, one drawback of reproducibility checklists when used for academic conferences and journals is the additional workload they impose on reviewers. To mitigate this workload, one option would be to leverage LLMs to assist the review process [44]. Finally, numerous guidelines and checklists for ML reproducibility have been recommended in various research fields [6]. Especially in Biomedical fields, we have seen a considerable adoption of guidelines and checklists, such as the TRIPOD statement [17], the CLAIM checklist [47], the ROBUST-ML checklist [2] or PROBAST [75].
@@
Model cards [46] are documentation sheets that provide information about ML models, including their intended use, potential limitations, and ethical considerations. They aim to enhance transparency in AI, by detailing aspects such as data used for training, performance metrics and evaluation methodologies, and possible biases. Model cards help users to understand and evaluate ML models more comprehensively, such that they are not being deployed in contexts other than the intended use, and thus to increase reproducibility. Similarly, model info sheets also provide documentation about ML models, but are specifically designed for the detection and prevention of data leakage in ML models [39]. The model info sheets are published alongside research to enable other researchers to quickly verify the validity of the data used to train ML models. Model info sheets require authors to answer detailed questions about the data and corresponding train/test splits, targeting various types of data leakage [39]. Despite that, they have two main drawbacks: first, verifying the correctness of info sheets only works after reproducing the results; second, completing these sheets requires a certain level of expertise in ML.
In general, model cards and model info sheets represent a promising, low-effort driver for ML reproducibility. They are especially useful in handling some of the methodological issues associated with ML models being developed by non-experts [39].
@@ 4.3 Awareness and Education
Awareness of reproducibility issues and available training/education to support reproducibility can be a very strong, contributing factor to improve the implemented practice (e.g., [74]).
@@
To enhance awareness and establish a minimum of reproducibility standards, the policies of scientific journals are considered an influence factor. A number of journals already mandate data and/or code availability for publication [57, 56, 30]. However, to address issues such as result manipulation, more extensive journal participation is needed to for instance introduce preregistration, where researchers register their research intentions for future publication. This approach ensures credibility by separating the research plan from experimental outcomes [70, 52], thereby reducing spin practices, HARKing, and p-hacking [29]. ACM TORS (Transactions on Recommender Systems) journal exemplifies this by allowing preregistration and publishing “reproducibility papers” dedicated to reproduction studies and enhancing reproducibility tools. Apart from that, various initiatives have been launched to raise awareness of reproducibility issues. A few examples are provided below:
@@
  • PapersWithCode.com [^19] https://paperswithcode.com/ is a resource for (1) ML papers, accompanied by the code, (2) datasets, and (3) ML methods. The ML papers include a link to a repository, which features the code and other artifacts for reproducing the results.
@@
  • The ACM has put together a new emerging interest group on reproducibility [^20] https://reproducibility.acm.org/. The main goals are to (1) contribute to the development of reproducibility standards, practices and policies, (2) promote the development and evaluation of tools, techniques, and methodologies and (3) encourage best practices.
  • ReproducedPapers.org is another online repository fostering reproductions. It further, focuses on education by incorporating a reproduction project into a Master’s level ML course at TU Delft [77].

### v2

@@ 4.1 Technology-based Drivers
Utilizing hosting services offers an efficient way to share code, data, and ML model parameter settings, thus supporting the reproducibility of ML-driven research [74]. Examples of hosting services include the runtime environments of ML platforms. If the original author runs the ML experiment in such a runtime environment, e.g., Kaggle Notebooks [^3] https://www.kaggle.com/code, Google Colab [^4] https://colab.google/, or CodaLab [^5] https://codalab.org/, researchers attempting to reproduce the results should be able to execute the experiment within the same environment.
@@
Reproducing the environment and setup of any ML experiment requires the consideration of existing dependencies and software versions, and is usually a complex task itself. Virtualization can simplify this process by bundling the essential components of ML models and experiments, such as the environment and code, into a single package for being shared with other researchers. Thus, if the authors of a paper build the experiment in a virtual environment, issues associated with setup reproduction can be greatly reduced. However, the adoption of virtualization by researchers depends on its user-friendliness and the effort of integration into their current workflows [12]. Concerns about virtualization include its limitations in allowing researchers to build upon them in a scalable manner. Traditional virtual machines (VMs) emulate an entire operating system for setting up and running experiments. The use of containerization software like Docker [^6] https://www.docker.com/ has just recently become popular. Containers are more lightweight and flexible than VMs, making it easier to adapt environments for follow-up studies [12]. There are also designated platforms for computational research, such as Code Ocean [^7] https://codeocean.com/, that offer virtualization via so-called reproducible capsules. Their focus in particular is to simplify the virtualization process and allow researchers to focus on the research questions instead of the standardization of environments [16]. Additionally, there are many other tools, such as e.g., ReproZip [15], which is one of the recommended tools by the SIGMOD Reproducibility [^8] https://reproducibility.sigmod.org/ to streamline reproducibility, and DetTrace [53], which aims to ensure completely deterministic computations.
@@
Many different sources of randomness during ML training lead to irreproducibility of ML research. Managing these sources of randomness, e.g., via random number seeds, deterministic algorithms, or other methods, could therefore greatly increase reproducibility. Fixed andom number seeds should be used and published to make ML experiments more reproducible and control a number of sources of the inherent nondeterminism. A seed is a first value used to initialize the pseudo-random number generator. When the same seed is used, the sequence of pseudo-random numbers generated is deterministic, meaning it will be the same every time the code is run. Additionally, ML models should be benchmarked and evaluated with multiple random number seeds, such that the variance can be reported and inform about the true performance of an ML model [65]. However, fixing the random number seeds can only eliminate algorithmic sources of randomness. Furthermore, hardware nondeterminism is also avoidable by using patching, i.e., replacing non-deterministic operations from hardware with deterministic ones, however, this will lead to a significant computing overhead [14]. To counteract inherent nondeterminism in reinforcement learning and achieve reproducible evaluations, there are frameworks, e.g. Gym-Ignition [22, 13], rl_reach [8], or MinAtar [81], which act as standardized benchmarking environments. Within them, different algorithms designed for the same task can be evaluated and compared against each other in a common environment. The frameworks counteract the effects of inherent nondeterminism by automatically controlling random seeds and evaluating algorithms over a number of runs. Additionally, it is an ongoing field of research to implement fully deterministic reinforcement learning algorithms [52] and make use of them within such frameworks [73].
@@
Privacy-preserving technologies support reproducibility, as they enable the collaborative training of ML models without sharing private or sensitive data. The main benefit of this is, that ML models can make use of larger, and more diverse data, thus helping to decrease bias and leading to more reproducible ML models. The main aim of Privacy-Preserving Machine Learning (PPML) is to facilitate the use of privacy-sensitive data to create better ML models, and, to allow data owners to collaboratively train ML models on private data. In that regard, PPML has several requirements. First, protecting the confidentiality of the training data. Second, preventing the leakage of sensitive information from ML model parameters and outputs, i.e., to hinder the re-identification of individuals. Third, achieving the listed security and privacy aspects while still preserving the utility of the ML model [79]. To achieve this, a number of different techniques are being used and developed, such as differential privacy, fully homomorphic encryption, secure multi-party computation, and federated learning [20]. These techniques are implemented in software libraries such as TensorFlow Privacy, PySyft, ML Privacy Meter, CryptFlow, or Crypten [7]. Furthermore, data can be made anonymous by removing identifiable personal information. However, if too much data is removed, the ML models may perform poorly. If not enough data is removed, it may still be possible to re-identify individuals by combining many different non-unique features [79]. An alternative approach to PPML is to generate synthetic data that captures the same information as the original data. A robust technique for creating such datasets can produce readily available datasets of nearly any size, as demonstrated in Biomedical fields. This approach has led to the development of Synthea, a software package designed to generate synthetic patient data and electronic health care records [76].
@@
There are many tools and platforms that assist in the implementation and management of ML models and ML-based applications. A recent study has evaluated 19 ML tools to gain insights into their concepts constituting reproducibility support [63]. As a result, five main pillars of ML reproducibility in tools and platforms were identified: (i) code versioning, (ii) data access, (iii) data versioning, (iv) experiment logging, and (v) pipeline creation. Most of these pillars are associated with managing and keeping track of different artifacts created during phases of the ML lifecycle (i.e., design, development, and deployment) as for instance, datasets, labels, code, logs, environment dependencies, random number seeds, or hyperparameters [69]. Each of these artifacts influences the final results of the ML model. Consequently, most tools aim to collect, store, and manage these artifacts, ensuring researchers can access and use them during reproduction attempts. Notable are also various tools and platforms for experiment tracking [69], such as:
  • DVC [^9] https://dvc.org/ is a version control system for ML projects with a command-line interface similar to Git [^10] https://git-scm.com/. It integrates with Git, supports cloud storage, and handles large versioning of datasets. DVC ensures full code and data provenance by enabling experiment tracking.
  • MLflow [^11] https://mlflow.org/ is an open-source tool for supporting ML experiment tracking, ML model deployment, and centralized model storage. Additionally, it provides an easy-to-use Web dashboard.
@@
  • dToolAI [32] collects and packages ML models together with supplemental information, such as hyperparameter settings, appropriate metadata, and persistent URIs for model training data. In contrast to the other tools, dToolAI is specifically tailored for deep learning models.
AutoML platforms, such as H2O Driverless AI [^12] https://h2o.ai, Google Cloud AutoML [^13] https://cloud.google.com/automl, DataRobot [^14] https://www.datarobot.com/platform) are a novel subcategory of ML tools that aim to aid with every aspect of the ML lifecycle, from data aggregation to model deployment. Thus, AutoML tools could facilitate more standardized ML models and also take care of tasks like hyperparameter optimization. It is however questionable how practical these tools are for reproducible ML research, since they often hide ML model optimization procedures. Sometimes, e.g., in the case of H2O Driverless AI, the features go a step further and promise to be able to prevent problems such as overfitting or target leakage [^15] https://h2o.ai/wiki/target-leakage//. In the case of target leakage, this is done by checking for a strong correlation between a feature and the target and then taking action, e.g., warning the user or automatically handling it. This is, however, a very simple solution to the problem and does not address the more complex cases of data leakage that are often present in research, e.g., temporal leakage.
@@ 4.2 Procedural Drivers
Due to a lack of shared datasets, currently, many researchers in ML-driven research - most notably in Biomedical fields - use individually acquired data [46]. The collection of such data is a time-consuming task and bears a significant risk of causing reproducibility issues, e.g., bias or data leakage. Often, the number of participants within the datasets is not very large and, thus, might not translate well to new contexts. Creating shared and standardized datasets can, therefore, (i) save researchers time in acquiring new data, (ii) facilitate the collaborative and independent maintenance and verification of data to minimize methodological errors, and (iii) support transferability and generalizability through the use of multi-institutional data [46]. In addition to the standardization of datasets, data cards [62] provide a consistent and comparable framework for reporting essential aspects of ML datasets. This includes information, e.g., about access restrictions, risks and limitations associated with the usage of the dataset, or any preprocessing steps, amongst many other contents, which are needed for reproducible ML development.
Another issue is the lack of standardized evaluation methods, which leads to reported performances of ML models often being overly optimistic [19]. To ensure the statistical significance of ML model evaluations, it is crucial to report performance as an aggregate of results obtained from multiple random runs, and with different random number seeds [17]. Furthermore, ML models should, if possible, be tested and evaluated on multiple different datasets [21]. This underscores the need for standardized evaluation methods, which can be supported by checklists or tools to prevent errors in this critical aspect of ML research. For this, similarly to data cards, model cards [48] are aimed to standardize the evaluation and reporting of the performance of ML models for a variety of use cases. Model cards should inform users about the possible applications of the ML model and its limitations. For the creation of the model cards, a toolkit is also being developed by Google [^16] https://research.google/blog/introducing-the-model-card-toolkit-for-easier-model-transparency-reporting/. In reinforcement learning, the creation of standardized evaluation pipelines is continually being researched to enable reproducible benchmarking of different reinforcement learning algorithms [41].
@@
There are many guidelines and checklists that outline best practices for increasing the reproducibility of ML. The guidelines are often aimed at specific parts of the ML workflow. For example, the FAIR guiding principles [^17] https://www.go-fair.org/fair-principles/ aim to improve the management and stewardship of scientific data by making scientific data (i) findable, (ii) accessible, (iii) interoperable, and (iv) reusable. Other guidelines promote the transparency and openness of scientific reporting in general, such as the TOP guidelines [^18] https://www.cos.io/initiatives/top-guidelines, which target journals. Similarly, checklists provide a simple framework for ensuring certain criteria are met. Checklists have been applied effectively in the past, e.g., in safety-critical systems, where they were used as early as in 1935 to complete pre-flight checks in Boeing airplanes. A promising example is the ML checklist proposed in [59], which has been suggested as best practice by researchers of different fields, e.g., in chemistry [6]. The checklist requests information about (i) the models and algorithms being used, (ii) theoretical claims in the research article, (iii) data, (iv) code, and (v) the ML experiment(s). However, one drawback of reproducibility checklists when used for academic conferences and journals is the additional workload they impose on reviewers. To mitigate this workload, one option would be to leverage LLMs to assist the review process [45]. Finally, numerous guidelines and checklists for ML reproducibility have been recommended in various research fields [6]. Especially in Biomedical fields, we have seen a considerable adoption of guidelines and checklists, such as the TRIPOD statement [18], the CLAIM checklist [49], the ROBUST-ML checklist [2], or PROBAST [78].
@@
Model cards [48] are documentation sheets that provide information about ML models, including their intended use, potential limitations, and ethical considerations. They aim to enhance transparency in AI, by detailing aspects such as data used for training, performance metrics, evaluation methodologies, and possible biases. Model cards help users to understand and evaluate ML models more comprehensively, such that they are not being deployed in contexts other than the intended use, and thus to increase reproducibility. Similarly, model info sheets also provide documentation about ML models, but are specifically designed for the detection and prevention of data leakage in ML models [40]. The model info sheets are published alongside research to enable other researchers to quickly verify the validity of the data used to train ML models. Model info sheets require authors to answer detailed questions about the data and corresponding train/test splits, targeting various types of data leakage [40]. Despite that, they have two main drawbacks: first, verifying the correctness of info sheets only works after reproducing the results; second, completing these sheets requires a certain level of expertise in ML. In general, model cards and model info sheets represent a promising, low-effort driver for ML reproducibility. They are especially useful in handling some of the methodological issues associated with ML models that could arise, e.g., when being developed by non-experts [40].
@@ 4.3 Awareness and Education
Awareness of reproducibility issues and available training/education to support reproducibility can be a powerful driver for ML reproducibility [77]).
@@
To enhance awareness and establish a minimum of reproducibility standards, the policies of scientific journals are considered an influence factor. A number of journals already mandate data and/or code availability for publication [59, 58, 31]. However, to address issues such as result manipulation, more extensive journal participation is needed to for instance introduce preregistration, where researchers register their research intentions for future publication. This approach ensures credibility by separating the research plan from experimental outcomes [72, 54], thereby reducing spin practices, HARKing, and p-hacking [30]. The ACM TORS (Transactions on Recommender Systems) journal exemplifies this by allowing preregistration and publishing “reproducibility papers” dedicated to reproduction studies and enhancing reproducibility tools. Apart from that, various initiatives have been launched to raise awareness of reproducibility issues. A few examples are the following:
@@
  • PapersWithCode.com [^20] https://paperswithcode.com/ is a resource for (i) ML papers, accompanied by the code, (ii) datasets, and (iii) ML methods. The ML papers include a link to a repository, which features the code and other artifacts for reproducing the results.
@@
  • The ACM has put together a new emerging interest group on reproducibility [^21] https://reproducibility.acm.org/. The main goals are to (i) contribute to the development of reproducibility standards, practices and policies, (ii) promote the development and evaluation of tools, techniques, and methodologies, and (iii) encourage best practices.
  • ReproducedPapers.org is another online repository fostering reproductions. It further focuses on education by incorporating a reproduction project into a Master’s level ML course at TU Delft [80].